OpenStudy (buttermequeasy):

I need help with a bionomial distribution. It has 4 parts. Please help. Six percent of people in a city have type O- blood. you randomly select 39 people in the city and ask them if their blood type is O-. Can the normal distribution be used to approximate the binomial distribution? A. Yes, because both np ≥ 5 and nq ≥ 5 B. No, because np < 5 C. No, because nq < 5 (a) find the probability that exactly 10 people say they have O- blood What is the indicated probability? (b)find the probability that at least 10 people say they have O- blood.

OpenStudy (buttermequeasy):

(c) find the probability that fewer than 10 people say they have O- blood. (d) a blood drive would to get at least five donors with O- blood. There are 100 donors. What is the probability that there will not be enough O- blood donors? Can the normal distribution be used to approximate the binomial distribution? A. No, because np < 5 B. No, because nq < 5 C. Yes, because both np ≥ 5 and nq ≥ 5 What is the probability that there will not be enough O- blood donors?

OpenStudy (anonymous):

so for the part where 39 people are randomly selected, what is n and what is p? once you know those, you can calculate n*p and since p+q = 1, q = 1-p and you can calculate to see if n*q >= 5

OpenStudy (buttermequeasy):

n= 39 p=0.06 q=.94 np= 2.34 nq= 36.66

OpenStudy (anonymous):

so you know how to answer that part, yes?

OpenStudy (anonymous):

which part is giving you trouble?

OpenStudy (buttermequeasy):

yes, the part giving me trouble is finding the probability

OpenStudy (buttermequeasy):

is B. No, because np < 5?

OpenStudy (anonymous):

good. to find the probability, you'll need a calculator. do you have one? if so, what kind?

OpenStudy (anonymous):

\[P(X=k)=\left(\begin{matrix}n \\ k\end{matrix}\right)\cdot p^k \cdot \left( 1-p \right)^{n-k}\]

OpenStudy (buttermequeasy):

i have a TI-84

OpenStudy (anonymous):

2nd VARS will get you to distributions, correct?

OpenStudy (buttermequeasy):

Yes

OpenStudy (anonymous):

from there, if you arrow down you'll find binompdf( and binomcdf( If you want to calculate the probability of a single value (like in part (a), P(X=10) ), use the binompdf( the inputs are n , p, x separated by commas. in order to calculate, hit 2nd, VARS, arrrow down to binompdf( then hit enter the hit 3, then 9 then , then . then 0 then 6 then , then 1 then 0 then enter.

OpenStudy (anonymous):

no comma after the 3, sorry should look like this: binompdf(39,.06,10)

OpenStudy (buttermequeasy):

okay, i got 6.39 x 10^-5 which i converted to 0.0000639 because i need to round to seven decimal places

OpenStudy (anonymous):

that's what i got too!

OpenStudy (buttermequeasy):

i also have to pick from a graph would the normal distribution not be used to approximate the binomial distribution?

OpenStudy (anonymous):

well, what did you get for the first part? remember that np < 5 so the normal isn't really appropriate. you should have a histogram with x approximately equal to np as the tallest.

OpenStudy (anonymous):

do you have pics of the graphs?

OpenStudy (anonymous):

sorry, i have to go. good luck!